# Game-Theoretic Defenses for Robust Conformal Prediction Against Adversarial Attacks in Medical Imaging

### v1

@@ 4.3. Addressing RQ3: Game-Theoretic Adversarial Attack
$\min_{f_{k},q_{k}}\max_{g_{j}}\sum_{i\in\mathcal{I}_{\textrm{test}}}\mathbb{E}\left[|\Gamma(x^{\prime}_{i})|\right],$
@@
where $x^{\prime}_{i}=g_{j}(x_{i},y_{i},f_{0})$ represents the adversarially perturbed input under attack $g_{j}$, and the non-conformity score associated with the perturbed input $s^{k}(x^{\prime}_{i},y_{i})$ is computed based on $f_{k}$, i.e., the outcome is determined by both parties. Here, $\Gamma(x^{\prime}_{i})$ is the prediction set constructed using the defensive model $f_{k}$ and the corresponding threshold $q_{k}$. As illustrated in Algorithm 3, $q_{k}$ is obtained as the maximum of all $q^{j}_{1-\alpha}$ of the scores computed using the defensive model $f_{k}$ on the calibration set attacked by $g_{j}$.
@@
This optimization framework seeks to identify the defensive model $f_{k}$ and threshold $q_{k}$ that offer the most robust defense against the worst-case attack $g_{j}$. By considering the defender and attacker within a game-theoretic paradigm, we systematically evaluate and enhance the resilience of conformal prediction sets against a diverse range of adversarial threats. This ensures that the coverage guarantees of conformal prediction remain intact even when subjected to strategic adversarial perturbations, thereby reinforcing the reliability and robustness of the predictive system in adversarial environments.
Procedure: The steps in Algorithm 4 involve a calibration and a evaluation set that are exchangeable with the test set. It works by applying each potential adversarial attack to the calibration set, training defensive models against each attack, computing quantile thresholds and corresponding prediction set sizes on the evaluation set to construct the payoff matrix. The zero-sum game is then solved to identify the optimal defensive strategy that minimizes the maximum prediction set size. Finally, the selected quantile threshold is used to construct conservative prediction sets for the test data.
@@
2:Optimal defensive strategy for constructing prediction sets $(\Gamma(x_{i}))_{x_{i}\in\mathcal{I}_{\textrm{test}}}$
@@
      4:for each attack function $g_{j},j=1,\dots,m$ do
@@
            Use $f_{k}(x^{\prime}_{i})$ to compute non-conformity scores $\{s^{j}(x^{\prime}_{i},y_{i})\}_{i\in\mathcal{I}_{\text{cal}}}$.
@@
            $q^{j}_{1-\alpha}$ as the $\lceil(1+|\mathcal{I}_{\text{cal}}|)(1-\alpha)\rceil$-th smallest score in $\{s^{j}(x^{\prime}_{i},y_{i})\}_{i\in\mathcal{I}_{\text{cal}}}$.
      5:Determine the maximum quantile threshold $q_{k}=\max_{j=1,\dots,m}q^{j}_{1-\alpha}$ for defensive model $f_{k}$.
@@
6:for each defensive model $f_{k},k=1,\dots,p$ do
      7:for each attack function $g_{j},j=1,\dots,m$ do
@@
            where $\Gamma(x^{\prime}_{i})=|\{y\in\mathcal{Y}\mid s^{j}(x^{\prime}_{i},y)\leq q_{k}\}|$, with $x^{\prime}_{i}$ the perturbed input attacked by $g_{j}$ and $s^{j}(x^{\prime}_{i},y)$ computed using $f_{k}$.
8:Construct the payoff matrix $P$ where each entry $P_{k,j}$ represents the expected prediction set size when the attacker selects attack $g_{j}$ and the defender selects defensive model $f_{k}$.
9:Solve the zero-sum game using the payoff matrix $P$ to identify the optimal defensive strategy.
10:Select defensive models $f_{k},\,k=1,\dots,p$ according to the optimal defensive strategies.
@@
11:for each test instance $x_{i},i\in\mathcal{I}_{\textrm{test}}$ do
@@ 5.1.1. Dataset Preparation
The experiments utilize the MedMNIST dataset suite, selected via the --dataset argument (e.g., PathMNIST). Three specific datasets are employed, each with unique characteristics. PathMNIST [45] comprises 97176 images of colon pathology categorized into nine classes, with each image sized at $28\times 28$ pixels. OrganAMNIST [45] consists of 34561 abdominal CT images distributed across eleven classes, also with a resolution of $28\times 28$ pixels. TissuMNIST [46, 45] contains 236386 images of human kidney cortex cells organized into eight categories. Each grayscale image in TissuMNIST is originally $32\times 32\times 7$ pixels, with 2D projections obtained by taking the maximum pixel value along the axial axis of each pixel and subsequently resized to $28\times 28$ grayscale images [47]. These datasets are preprocessed to ensure uniformity across models, including normalization and resizing as required by the ResNet18 architecture.
@@ 5.1.2. Data Splitting and Experimental Setup
For each dataset and every defensive model, including the normal (non-adversarial) model, the data is split into training, validation, and test sets. Specifically, 50% of the training dataset is used to train the defensive models, while 10% is allocated for validation purposes. The remaining data constitutes the test set, which is further divided based on the research question being addressed. For RQ1 and RQ2, the test dataset is split equally into a calibration set and a test set, each comprising 50% of the original test data. In contrast, for RQ3, the test dataset is partitioned into three subsets: 25% for calibration, 25% for evaluation, and 50% for the final test set. This stratified splitting ensures that each class is appropriately utilized for training, validating, and evaluating the defensive models under different adversarial attack scenarios.
@@ 5.1.3. Adversarial Attack Types
Adversarial attacks are implemented using TorchAttacks [48]. The selected attacks include FGSM, PGD, and SPSA. Each attack type is designed to target the defensive model against which it is trained, facilitating comprehensive evaluation of model robustness across a variety of adversarial strategies. Employing multiple attack methodologies allows for a thorough assessment of a model’s vulnerabilities and ensures that defenses are not tailored to only one specific type of adversarial manipulation.
@@ 5.2. Results
5.2.1. Results for RQ1
In the first experiment, an attack of a known type is applied to the test data. We train defensive models $f_{j}$ against each attack $g_{j}$, for $j=1,\dots,m$, and a normal model $f_{0}$. We evaluate both the prediction accuracy and the conformal prediction set’s performance in terms of coverage, size, and SSCV. Specifically, Figure 1 illustrates the confusion matrices for both prediction accuracy and prediction set size for each defensive model and each attack. Comprehensive performance results are presented in Tables 2, 3, and 4. The key findings from RQ1 are as follows:
Defensive Model Accuracy
Although intuitively, the defensive model adversarially trained against a specific attack should exhibit the highest accuracy against that attack, this is not always the case. Various factors could be the reason, including the inherent complexity of certain attacks, the capacity of the defensive models, and the diversity of adversarial perturbations. Consequently, some defensive models demonstrate superior performance even against attacks they were not explicitly trained on.
@@ 5.2.1. Results for RQ1
The model with the best accuracy does not necessarily produce the most efficient conformal prediction sets. This discrepancy may be influenced by the chosen non-conformity score function. In our experiments, we employed the APS (Adaptive Prediction Set) [51] as the score function, which considers the cumulative probability of labels that have the same or lower estimated probabilities compared to the label of interest. As a result, APS emphasizes the tail probabilities, potentially leading to larger prediction sets for models focused primarily on accuracy.
(a) Accuracy
(b) Size
(c) Accuracy
(d) Size
(e) Accuracy
(f) Size
Figure 1. Confusion Matrices for Accuracy and Size across Datasets. Each row represents a dataset ($\bf{PathMNIST}$, $\bf{OrganAMNIST}$, $\bf{TissueMNIST}$) and each column represents a metric (Accuracy, Size).

### v2

@@ 4.3. Addressing RQ3: Game-Theoretic Adversarial Attack
$\min_{f_{k},q_{1-\alpha}^{k}}\max_{g_{j}}\sum_{i\in\mathcal{I}_{\textrm{test}}}\mathbb{E}\left[|\Gamma(x^{\prime}_{i})|\right],$
@@
where $x^{\prime}_{i}=g_{j}(x_{i},y_{i},f_{0})$ represents the adversarially perturbed input under attack $g_{j}$, and the non-conformity score associated with the perturbed input $s^{k}(x^{\prime}_{i},y_{i})$ is computed based on $f_{k}$, i.e., the outcome is determined by both parties. Here, $\Gamma(x^{\prime}_{i})$ is the prediction set constructed using the defensive model $f_{k}$ and the corresponding threshold $q_{k}$. As illustrated in Algorithm 3, $q_{k}$ is obtained as the maximum of all $q^{j}_{1-\alpha}$ of the scores computed using the defensive model $f_{k}$ on the calibration set attacked by $g_{j}$. To rigorously support our claim that the coverage guarantee is maintained even in the presence of strategic adversarial perturbations, we state and prove the following theorem.
Theorem 4.4 (Robust Coverage Guarantee).
Let $\{f_{k}\}_{k=1}^{p}$ be a set of predefined defensive models and $\{g_{j}\}_{j=1}^{m}$ be a set of adversarial attacks. Suppose that:
  1. (1)
    The calibration set $\mathcal{I}_{\text{cal}}$ and the test $\mathcal{I}_{\text{test}}$ set are exchangeable.
  2. (2)
    For any given defensive model $f_{k}$ and any attack $g_{j}$, the nonconformity score $s^{k}(x^{\prime}_{i},y_{i})$ satisfies regularity conditions that ensure the existence and proper ordering of quantiles (e.g., measurability and monotonicity properties).
  3. (3)
    The threshold $q_{1-\alpha}^{k}$ is defined as
    $q_{1-\alpha}^{k}=\max_{j\in\{1,\ldots,m\}}q^{k_{j}}_{1-\alpha},$
    where each $q^{k_{j}}_{1-\alpha}$ is computed on $\mathcal{I}_{\text{cal}}$ so that
    $\frac{1}{|\mathcal{I}_{\text{cal}}|}\sum_{i\in\mathcal{I}_{\text{cal}}}\mathbf{1}\{s_{k}^{j}(x^{\prime}_{i},y_{i})>q^{k_{j}}_{1-\alpha}\}\leq\alpha.$
Then, for every attack $g_{j}$, the prediction set
$\Gamma(x^{\prime}_{i})=\{y\in\mathcal{Y}:s_{k}^{j}(x^{\prime}_{i},y)\leq q_{1-\alpha}^{k}\}$
satisfies the coverage guarantee
$\mathbb{P}(y\in\Gamma(x^{\prime}_{i}))\geq 1-\alpha.$
Proof.
Fix any defensive model $f_{k}$ and an arbitrary attack $g_{j}$. By the construction on the calibration set under the attack $g_{j}$, the quantile $q^{k_{j}}_{1-\alpha}$ satisfies
$\frac{1}{|\mathcal{I}_{\text{cal}}|}\sum_{i\in\mathcal{I}_{\text{cal}}}\mathbf{1}\Big{\{}s_{k}^{j}(x^{\prime}_{i},y_{i})>q^{k_{j}}_{1-\alpha}\Big{\}}\leq\alpha.$
This implies
$\mathbb{P}\Big{(}s_{k}^{j}(x^{\prime}_{i},y_{i})\leq q^{k_{j}}_{1-\alpha}\Big{)}\geq 1-\alpha,$
where the probability is taken with respect to the data distribution (leveraging the exchangeability or i.i.d. assumption between the calibration and test sets).
Since the threshold $q^{k}_{1-\alpha}$ is defined as the maximum over all attacks, we have
$q^{k}_{1-\alpha}=\max_{1\leq j\leq m}q^{k_{j}}_{1-\alpha}\geq q^{k_{j}}_{1-\alpha}.$
Under the assumption of the monotonicity of the indicator function, i.e.,
$\{s^{k}(x^{\prime}_{i},y_{i})\leq q^{k_{j}}_{1-\alpha}\}\subseteq\{s^{k}(x^{\prime}_{i},y_{i})\leq q^{k}_{1-\alpha}\},$
it directly follows that
$\mathbb{P}\Big{(}s^{k}(x^{\prime}_{i},y_{i})\leq q_{k}\Big{)}\geq\mathbb{P}\Big{(}s^{k}(x^{\prime}_{i},y_{i})\leq q^{j_{0}}_{1-\alpha}\Big{)}\geq 1-\alpha.$
Since $j$ was arbitrary, this argument holds for any attack $g_{j}$. Hence, the prediction set constructed with $q^{k}_{1-\alpha}$ satisfies the desired coverage guarantee:
$\mathbb{P}(y\in\Gamma(x^{\prime}_{i}))\geq 1-\alpha.$
@@
This optimization framework seeks to identify the defensive model $f_{k}$ and threshold $q_{k}$ that offer the most robust defense against the worst-case attack $g_{j}$. It further weighs the balance between coverage rate and the size of conformal prediction sets. By framing the interaction between the defender and attacker within a game-theoretic context, the framework systematically assesses and bolsters the resilience of conformal prediction sets against a multitude of adversarial threats. Simultaneously, the defender strives to minimize the size of the prediction set to decrease model uncertainty, thereby ensuring that the predictive system maintains coverage guarantees while offering a smaller, more credible range for medical imaging decisions, as excessively large prediction sets lack practical significance.
Procedure: The steps outlined in Algorithm 4 involve utilizing a calibration set and an evaluation set, which are interchangeable with the test set. The process begins by applying each potential adversarial attack to the calibration set. Subsequently, defensive models are trained against each attack. Quantile thresholds and corresponding prediction set sizes are then computed on the evaluation set to construct the payoff matrix. A key aspect of this algorithm, specifically in Step 4, is the utilization of the NashPy package to solve the zero-sum game. This involves employing an enumerative approach to iterate through all possible support set combinations, constructing and solving the corresponding linear equation systems to identify mixed strategies that satisfy the Nash equilibrium conditions. For each identified strategy combination, NashPy verifies its non-negativity and normalization, ultimately returning all feasible mixed strategies as the Nash equilibria of the game. Finally, the selected quantile threshold is applied to construct conservative prediction sets for the test data.
@@
2:Optimal defensive strategy for constructing prediction sets $\left(\Gamma(x_{i})\right)_{x_{i}\in\mathcal{I}_{\textrm{test}}}$
@@
      for each attack function $g_{j},j=1,\dots,m$ do
@@
            Use $f_{k}(x^{\prime}_{i})$ to compute non-conformity scores $\{s_{k}^{j}(x^{\prime}_{i},y_{i})\}_{i\in\mathcal{I}_{\text{cal}}}$.
@@
            $q^{k_{j}}_{1-\alpha}$ as the $\lceil(1+|\mathcal{I}_{\text{cal}}|)(1-\alpha)\rceil$-th smallest score in $\{s_{k}^{j}(x^{\prime}_{i},y_{i})\}_{i\in\mathcal{I}_{\text{cal}}}$.
      Determine the maximum quantile threshold $q_{1-\alpha}^{k}=\max_{j=1,\dots,m}q^{k_{j}}_{1-\alpha}$ for defensive model $f_{k}$.
@@
4:for each defensive model $f_{k},k=1,\dots,p$ do
      for each attack function $g_{j},j=1,\dots,m$ do
@@
            where $\Gamma(x^{\prime}_{i})=|\{y\in\mathcal{Y}\mid s_{k}^{j}(x^{\prime}_{i},y)\leq q_{1-\alpha}^{k}\}|$, with $x^{\prime}_{i}$ the perturbed input attacked by $g_{j}$ and $s_{k}^{j}(x^{\prime}_{i},y)$ computed using $f_{k}$.
$\triangleright$ Optimal Defensive Strategy Derivation under Nash Equilibrium:
5:Construct the payoff matrix $P$ where each entry $P_{k,j}$ represents the expected prediction set size when the attacker selects attack $g_{j}$ and the defender selects defensive model $f_{k}$.
6:Solve the zero-sum game using the payoff matrix $P$ to identify the optimal defensive strategy.
7:Select defensive models $f_{k},\,k=1,\dots,p$ according to the optimal defensive strategies.
@@
8:for each test instance $x_{i},i\in\mathcal{I}_{\textrm{test}}$ do
@@ 5.1.1. Dataset Preparation
The experiments utilize the MedMNIST dataset suite [49, 50]. Three specific datasets are employed, each with unique characteristics. The datasets are selected via the --dataset command-line argument, which allows the user to specify one of the available datasets (e.g., PathMNIST, OrganAMNIST, or TissuMNIST). This design provides flexibility in choosing different organ-specific visual recognition tasks and facilitates reproducibility of the experiments.
Specifically, PathMNIST comprises 97,176 images of colon pathology, each of size $28\times 28$ pixels, categorized into nine classes. OrganAMNIST contains 34,561 abdominal CT scans, similarly resized to $28\times 28$ pixels, and distributed across eleven classes. TissuMNIST includes 236,386 images of human kidney cortex cells originally sized $32\times 32\times 7$. For TissuMNIST, 2D projections are obtained by taking a maximum pixel value along the axial axis, and the resulting images are resized to $28\times 28$ pixels. For the predictive models, all experiments deploy the ResNet18 architecture as the backbone [51].
Overall, the experimental setup is designed to rigorously evaluate both standard and adversarially-trained models across multiple datasets and attack scenarios, providing insights into model performance and adversarial resilience.
@@ 5.1.2. Data Splitting and Experimental Setup
For each dataset and every defensive model, including the normal (non-adversarial) model, we adopt a stratified splitting procedure to divide the data into training, validation, and test sets. This approach is designed to preserve the original class distribution across all splits. Specifically, 50% of the training data is used to train the defensive models, while 10% is allocated for validation purposes. The remaining data constitutes the test set.
The test set is then further subdivided based on the research question to ensure balanced representation:
  • RQ1 and RQ2: The test dataset is split equally into a calibration set and a test set, each comprising 50% of the original test data.
  • RQ3: The test dataset is partitioned into three subsets: 25% for calibration, 25% for evaluation, and 50% for the final test set.
All experimental settings use an alpha value of 0.1. This stratified splitting procedure ensures that each class is appropriately utilized for training, validation, and evaluation, thereby supporting consistent performance assessment of the defensive models under different adversarial attack scenarios.
@@ 5.1.3. Adversarial Attack Types
Adversarial attacks are implemented using TorchAttacks [52]. The selected attacks include FGSM, PGD, SPSA, CW and Clean. Each attack type is designed to target the defensive model against which it is trained, facilitating comprehensive evaluation of model robustness across a variety of adversarial strategies. Employing multiple attack methodologies allows for a thorough assessment of a model’s vulnerabilities and ensures that defenses are not tailored to only one specific type of adversarial manipulation.
@@ 5.2. Results
The model with the best accuracy does not necessarily produce the most efficient conformal prediction sets. This discrepancy may be influenced by the chosen non-conformity score function. In our experiments, we evaluated four conformal prediction (CP) methods: the standard APS [55], RSCP [31], VRCP-I [38], and VRCP-C [38]. The standard APS method uses the cumulative probability of labels that have the same or lower estimated probability than the target label, emphasizing the tail probabilities. RSCP extends this approach by incorporating Gaussian noise into the inputs and averaging the scores across multiple samples, effectively “smoothing” the estimated scores. In contrast, VRCP-I and VRCP-C introduce a robustness component by applying adversarial perturbations to the inputs within a specified $\epsilon$-ball. Specifically, VRCP-I computes the minimum score over multiple perturbed copies (thus capturing a more conservative score for each class), while VRCP-C focuses on the worst-case score across these perturbations. These methods collectively allow us to dynamically adjust prediction sets to better capture uncertainty, with the RSCP and VRCP variants enhancing robustness under adversarial conditions, which may result in larger prediction sets for accuracy-driven models.
5.2.1. Results for RQ1 and RQ2
In our experimental evaluation, we compared two distinct scenarios: one in which the attack’s nature is known (RQ1), and another in which we assume that an unknown attack belongs to one of five predefined categories (FGSM, PGD, SPSA, CW, or Clean) (RQ2). Both experiments are conducted under the same conformal prediction (CP) framework, which enables us to directly assess how the CP coverage and prediction set sizes vary under different conditions. Additionally, in Tables 3 to 6, the bold entries highlight the best defense model for a particular CP method under each type of attack, while underlined entries indicate the corresponding metric for the best CP method under a specific attack and defense model combination.
For RQ1, known attacks are applied to the test data. For each attack $g_{j}$ (where $j=1,\dots,m$), we train a corresponding defensive model $f_{j}$ alongside a baseline model $f_{0}$. In contrast, for RQ2 we assume that the unknown attack belongs to one of the five designated categories and apply it to the test data. We then employ Algorithm 3 to determine the maximum threshold obtained across the calibration sets under different attacks, thereby facilitating the construction of a conservative CP that guarantees the required coverage level. To better highlight the connections and distinctions between RQ1 and RQ2, we discuss the results from both scenarios together.